# BalancEdit: Dynamically Balancing the Generality-Locality Trade-off in Multi-modal Model Editing

### v1

@@ Abstract
Large multi-modal models inevitably decay over time as facts change and previously learned information becomes outdated. Traditional approaches such as fine-tuning are often impractical for updating these models due to their size and complexity. Instead, direct knowledge editing within the models presents a more viable solution. Current model editing techniques, however, typically overlook the unique influence ranges of different facts, leading to compromised model performance in terms of both generality and locality. To address this issue, we introduce the concept of the generality-locality trade-off in multi-modal model editing. We develop a new model editing dataset named OKEDIT, specifically designed to effectively evaluate this trade-off. Building on this foundation, we propose BalancEdit, a novel method for balanced model editing that dynamically achieves an optimal balance between generality and locality. BalancEdit utilizes a unique mechanism that generates both positive and negative samples for each fact to accurately determine its influence scope and incorporates these insights into the model’s latent space using a discrete, localized codebook of edits, without modifying the underlying model weights. To our knowledge, this is the first approach explicitly addressing the generality-locality trade-off in multi-modal model editing. Our comprehensive results confirm the effectiveness of BalancEdit, demonstrating minimal trade-offs while maintaining robust editing capabilities. Our code and dataset will be available.
@@ 1 Introduction
Large multi-modal models (Zhu et al., 2023; Radford et al., 2021; Li et al., 2023; Liu et al., 2023a; Rombach et al., 2022) have recently brought about significant advancements in artificial intelligence, demonstrating impressive results in tasks such as Visual Question Answering (VQA) (Antol et al., 2015). However, these models are susceptible to issues like hallucination (Rawte et al., 2023) and fact alteration (De Cao et al., 2021). After deployment, these models may generate numerous errors, leading to potential problems like the propagation of hate speech or the dissemination of outdated factual information. Given these challenges, it is critical to continually update and maintain these large multi-modal models to ensure their accuracy and relevance.
While retraining or fine-tuning can update a model’s knowledge, it is often infeasible to frequently edit individual facts due to the high computational costs involved. Fortunately, model editing techniques (Hartvigsen et al., 2024; Mitchell et al., 2021; Zheng et al., 2023) provide a promising approach to implementing cost-effective, targeted updates to large, pre-trained models. These techniques typically involve injecting new layers or modifying weights to alter the knowledge embedded in language models. A successful edit generally exhibits three characteristics (Mitchell et al., 2021; Huang et al., 2023): reliability, which ensures the output changes to the target answer for the same question; locality, which leaves unrelated knowledge and outputs unchanged; and generality, which produces the correct answer for all questions within the influence scope. As illustrated in Fig. 1, each fact has its own influence scope. For instance, if we wish to edit the name of a specific cat, the influence scope would be confined to that particular cat. If we aim to edit the name of a cat breed, the influence scope would extend to all cats within that breed. However, if we intend to edit the name of a species, the influence scope would encompass all cats. Consequently, we should consider each fact individually and dynamically to determine the appropriate influence scope.
@@
However, current model editing techniques often overlook the dynamic nature of the influence scope. Some methods treat all influence scopes as if they are large and uniform, while others focus solely on a specific edit. For instance, IKE (Zheng et al., 2023) employs in-context learning to edit knowledge, using the closest piece of knowledge as a prompt to guide the language model. This approach causes the language model to rephrase the nearest fact, resulting in an oversized influence scope. Conversely, GRACE (Hartvigsen et al., 2024), a lifelong model editing method, assumes each edit has a small and similar influence range, leading to limited generality. Consider an example where we aim to edit a “fact” that HP computers have been renamed Lenovo, as shown in Table 1. Ideally, model editing should update the answer from HP to Lenovo whenever it encounters an image of a HP computer, while leaving the answer unchanged for other brands. However, existing model editing techniques, such as IKE (Zheng et al., 2023) and MEND (Mitchell et al., 2021), may achieve the target edit but neglect the influence scope, inadvertently editing other brands as well. Even when presented with a black image, these models may still output the new answer, leading to hallucination. On the other hand, while GRACE (Hartvigsen et al., 2024) maintains the backbone model’s answer for unrelated images, it fails to edit the knowledge to the desired scope. These observations suggest that existing multi-modal model editing methods struggle to dynamically adjust the influence scope of a knowledge edit, and to balance generality and locality effectively
@@
To address this issue, we first create a dataset designed to evaluate the trade-off between generality and locality in model editing techniques. We then introduce an efficient multi-modal model editing method named BalancEdit, which dynamically balances this trade-off with minimal computational costs. Specifically, we incorporate an adapter into a chosen layer of a vision language model without altering its weights. This adaptor modifies layer-to-layer transformations for select inputs. By caching embeddings for input errors and the updated knowledge transformation layer that decodes into the desired model outputs, BalancEdit functions as a codebook where edits are stored. To strike a balance between generality and locality, we generate the corresponding positive and negative samples for each edit. The model’s semantic similarity in its latent space can be visualized as dynamic spheres around cached edits, with the radius determined by the distance between positive and negative samples. By adjusting the radius over time, BalancEdit allows for immediate edits, retains previous edits, and preserves correct model behaviors, making it parameter-efficient. Furthermore, since BalancEdit’s codebooks do not alter model weights and are fully model-agnostic, they also pave the way for plug-and-play, cost-effective model editing. This is particularly useful for making critical spot-fixes between larger retraining efforts.
@@ 2 Related Work
Large Vision Language Models. Vision language models (Radford et al., 2021; Zhu et al., 2023; Li et al., 2023, 2022; Wang et al., 2024; Zhou et al., 2024; Lin et al., 2024; Dai et al., 2024) are one of the key part in multi-modal learning, which aim to learn multi-modal foundation models with improved performance on vision language tasks, such as VQA (Antol et al., 2015). These models (Li et al., 2022; Liu et al., 2023a), by mapping image embeddings to text embedding space, are capable of interpreting image information and handling a wide array of tasks. They demonstrate impressive abilities in image understanding, generation, and reasoning. These capabilities, however, rely heavily on millions of high-quality training data (Schuhmann et al., 2022, 2021). Given that factual knowledge, especially visual information, changes over time, it is crucial to keep the model up-to-date. However, updating the model’s behavior through retraining or fine-tuning is impractical due to exorbitant training costs. In this context, multi-modal model editing techniques, which allow for targeted edits, provide a feasible solution to this challenge.
@@ 3.2 BalancEdit
  • Keys ($K$): Each key $\bm{k}$ stores the averaged embedding produced by the layer $l-1$ for a specific question answer pair. Mathematically, it can be expressed as $K=\{\bm{k}=\bar{\bm{h}}^{l-1}_{i,t}|\bar{\bm{h}}^{l-1}_{i,t}=\frac{1}{n}\sum f^{l-1}(\bm{i,t}),\forall\bm{(i,t)}\in D_{\text{edit}}\}$.
  • Transformations ($V$): Each transformation $v(\cdot)$ associated with a specific key $\bm{k}$ stores the new weights with the updated knowledge. Typically, the transformation is fine-tuned with the model’s finetuning loss with updated knowledge.
@@
$v^{*}=\arg\min_{v}\bm{L}(f_{new}(i,t),y^{n})$ (2)
Specifically, if the key is empty or the new fact falls outside the influence scope of existing keys, the transformation is directly finetuned from the original transformation layer. However, there may be instances where the new fact overlaps with the existing keys. In such cases, we finetune the transformation layer from the previously edited transformation to prevent catastrophic forgetting. Additionally, if the new key directly conflicts with previous edits, we will discard the previous entry and add a new one to update the knowledge. To ensure the universality, we primarily utilize the basic full fine-tuning approach as the transformation method. This involves adjusting the weights of the neural network to better align with the newly introduced or modified knowledge without altering the overall architecture of the model. The parameters that are tuned include all the weights within the specific layer of the network.
@@ 4.1 Datasets and Backbone Models
Backbone Models. Following previous work (Zheng et al., 2023), we adopt two vision language models as the base models. MiniGPT-4 (Zhu et al., 2023) is a powerful vision language model, leveraging Vicunna (Chiang et al., 2023) as the language model and a Vit-G/14 from EVA-CLIP (Sun et al., 2023) and a Q-former as the image encoder. BLIP-2 OPT (Li et al., 2023) utilizes a lightweight Q-former to bridge the gap between vision modality and text modality, where the ViT-L is adopted in the vision block, and the unsupervised-trained OPT model (Zhang et al., 2022) is used for decoder-based LLM. Metrics Following previous work (Zheng et al., 2023), we adopt the Editing Success Accuracy (Acc); Text Generality (T-Gen); Image Generality (I-Gen); and Locality (Loc) as the main metrics. To quantify the trade-off between generality and locality, we introduce the harmonic mean (HM) of the T-Gen, I-Gen and Loc. The detailed informations are in Appendix C.
4.2 Baselines
4.3 Comparisons to Existing Methods
@@ 4.5 Efficiency Evaluation
Table 5: Time efficiency evaluation results on BLIP2-OPT.
Time Efficiency. Training time is divided into two components, as detailed in Table 5. The first component is pre-training time, which involves either pre-training the model editing method or preparing the augmented index. For instance, MEND, a meta-learning method, requires 22 hours to pre-train on 6,346 training samples. IKE, a retrieval-augmented in-context learning method, needs 12 hours to index 6,346 knowledge facts in advance.
@@ 4.7 Ablation Study
In the ablation study, we test the influences on the hyperparameter $\alpha$ to show the trade-off between generality and locality. In addition to that, we compare the different distance functions to show the generalization ability of our BalancEdit.
@@
Effect of the Hyperparameter. In this study, we conducted a series of experiments on a subset of the OKVQA dataset to investigate how the parameter $\alpha$ affects the trade-off between generality and locality in model editing. As illustrated in Figure 5, we varied $\alpha$ from 0.1 and 0.3. The results, depicted in the figure, show that the editing success accuracy and text generality metrics consistently maintain a 100% accuracy rate. This stability is attributed to these metrics being closely tied to the key, with changes in the radius having no significant impact on them. However, the image generality metric, which is more challenging, shows a decline as $\alpha$ increases. This trend is anticipated because questions related to image generality tend to deviate from the key, despite sharing similar semantic content. Consequently, as the radius decreases, the edited model tends to overlook these questions. Conversely, the model’s performance on locality improves with an increase in $\alpha$. A smaller radius helps preserve the integrity of unrelated questions, ensuring that their answers remain unchanged. In this scenario, we observe that the harmonic mean of generality and locality initially increases and then decreases, further validating the existence of this trade-off. However, our method continues to achieve relatively high performance.
@@
Effect of Alternative Negative Anchors: To further validate the effectiveness of our approach, we conducted experiments using various negative anchors, including white negative images, on a subset of the OKEDIT dataset. As shown in Table 7, both black and white negative samples achieved 100% editing accuracy and exhibited a high harmonic mean in the locality-generality trade-off. The performance metrics for both white and black negative anchors, such as accuracy, generalization metrics, and locality, are remarkably consistent. The slight variations in the locality and I-Gen metrics suggest that white images can function as effective negative anchors, which also lack significant discriminative information. This consistency across different negative anchors highlights the robustness and adaptability of our pipeline in various settings and confirms our assumptions.

### v2

@@ Abstract
Large multi-modal models inevitably decay over time as facts update and previously learned information becomes outdated. Traditional approaches such as fine-tuning are often impractical for updating these models due to their size and complexity. Instead, direct knowledge editing within the models presents a more viable solution. Current model editing techniques, however, typically overlook the unique influence ranges of different facts, leading to compromised model performance in terms of both generality and locality. To address this issue, we introduce the concept of the generality-locality trade-off in multi-modal model editing. We develop a new model editing dataset named OKEDIT, specifically designed to effectively evaluate this trade-off. Building on this foundation, we propose BalancEdit, a novel method for balanced model editing that dynamically achieves an optimal balance between generality and locality. BalancEdit utilizes a unique mechanism that generates both positive and negative samples for each fact to accurately determine its influence scope and incorporates these insights into the model’s latent space using a discrete, localized codebook of edits, without modifying the underlying model weights. To our knowledge, this is the first approach explicitly addressing the generality-locality trade-off in multi-modal model editing. Our comprehensive results confirm the effectiveness of BalancEdit, demonstrating minimal trade-offs while maintaining robust editing capabilities. Our code and dataset are available at https://github.com/donglgcn/BalancEdit/tree/MMOKVQA.
@@ 1 Introduction
Large multi-modal models (Zhu et al., 2023; Radford et al., 2021; Li et al., 2023; Liu et al., 2023a; Rombach et al., 2022) have recently brought about significant advancements in artificial intelligence, demonstrating impressive results in tasks such as Visual Question Answering (VQA) (Antol et al., 2015). However, these models are susceptible to issues like hallucination (Rawte et al., 2023) and fact alteration (De Cao et al., 2021; Guo et al., 2024; Zhu et al., 2024). After deployment, these models may generate numerous errors, leading to potential problems like the propagation of hate speech (Guan et al., 2025b, a; Hu et al., 2025, 2024) or the dissemination of outdated factual information. Given these challenges, it is critical to continually update and maintain these large multi-modal models to ensure their accuracy and relevance.
While retraining or fine-tuning can update a model’s knowledge, it is often infeasible to frequently edit individual facts due to the high computational costs involved. Fortunately, model editing techniques (Hartvigsen et al., 2024; Mitchell et al., 2021; Zheng et al., 2023) provide a promising approach to implementing cost-effective, targeted updates to large pretrained models. These techniques typically involve injecting new layers or modifying weights to alter the knowledge embedded in language models. A successful edit generally exhibits three characteristics (Mitchell et al., 2021; Huang et al., 2023): reliability, which ensures the output changes to the target answer for the same question; locality, which leaves unrelated knowledge and outputs unchanged; and generality, which produces the correct answer for all questions within the influence scope. As illustrated in Fig. 1, each fact has its own influence scope. For instance, if we wish to edit the name of a specific cat, the influence scope would be confined to that particular cat. If we aim to edit the name of a cat breed, the influence scope would extend to all cats within that breed. However, if we intend to edit the name of a species, the influence scope would encompass all cats. Consequently, we should consider each fact individually and dynamically to determine the appropriate influence scope.
@@
However, current model editing techniques often overlook the dynamic nature of the influence scope. Some methods treat all influence scopes as if they are large and uniform, while others focus solely on a specific edit. For instance, IKE (Zheng et al., 2023) employs in-context learning to edit knowledge, using the closest piece of knowledge as a prompt to guide the language model. This approach causes the language model to rephrase the nearest fact, resulting in an oversized influence scope. Conversely, GRACE (Hartvigsen et al., 2024), a lifelong model editing method, assumes that each edit has a small and similar influence range, leading to limited generality. Consider an example where we aim to edit a “fact” that HP computers have been renamed Lenovo, as shown in Table 1. Ideally, model editing should update the answer from HP to Lenovo whenever it encounters an image of a HP computer, while leaving the answer unchanged for other brands. However, existing model editing techniques, such as IKE (Zheng et al., 2023) and MEND (Mitchell et al., 2021), may achieve the target edit but neglect the influence scope, inadvertently editing other brands as well. Even when presented with a black image, these models may still output the new answer, leading to hallucination. On the other hand, while GRACE (Hartvigsen et al., 2024) maintains the backbone model’s answer for unrelated images, it fails to edit the knowledge to the desired scope. These observations suggest that existing multi-modal model editing methods struggle to dynamically adjust the influence scope of a knowledge edit, and to balance generality and locality effectively.
@@
To address this issue, we first create a dataset designed to evaluate the trade-off between generality and locality in model editing techniques. We then introduce an efficient multi-modal model editing method named BalancEdit, which dynamically balances this trade-off with minimal computational costs. Specifically, we incorporate an adapter into a chosen layer of a vision language model without altering its weights. This adaptor modifies layer-to-layer transformations for select inputs. By caching embeddings for input edits and the updated knowledge transformation layer, BalancEdit functions as a codebook where edits are stored. To strike a balance between generality and locality, we generate the corresponding positive and negative samples for each edit. The model’s semantic similarity in its latent space can be visualized as dynamic spheres around cached edits, with the radius determined by the distance between positive and negative samples. By adjusting the radius over time, BalancEdit allows for immediate edits, retains previous edits, and preserves correct model behaviors. Furthermore, since BalancEdit’s codebooks do not alter model weights and are fully model-agnostic, they also pave the way for plug-and-play, cost-effective model editing. This is particularly useful for making critical spot-fixes between larger retraining efforts.
@@ 2 Related Work
Large Vision Language Models. Vision language models (Radford et al., 2021; Zhu et al., 2023; Li et al., 2023, 2022; Wang et al., 2024a; Zhou et al., 2024; Lin et al., 2024; Dai et al., 2024) are one of the key part in multi-modal learning, which aim to learn multi-modal foundation models with improved performance on vision language tasks (Antol et al., 2015; Guo et al., ; Wang et al., 2024b). These models (Li et al., 2022; Liu et al., 2023a), by mapping image embeddings to text embedding space, are capable of interpreting image information and handling a wide array of tasks. They demonstrate impressive abilities in image understanding, generation, and reasoning. These capabilities, however, rely heavily on millions of high-quality training data (Schuhmann et al., 2022, 2021). Given that factual knowledge, especially visual information, changes over time, it is crucial to keep the model up-to-date. However, updating the model’s behavior through retraining or fine-tuning is impractical due to exorbitant training costs. In this context, multi-modal model editing techniques, which allow for targeted edits, provide a feasible solution to this challenge.
@@ 3.2 BalancEdit
  • Keys ($K$): Each $k$ in the codebook is a representative embedding for a specific edit. Each $\bm{k}$ stores the averaged embedding produced by the layer $l-1$ for a specific question answer pair. Mathematically, it can be expressed as $K=\{\bm{k}=\bar{\bm{h}}^{l-1}_{i,t}|\bar{\bm{h}}^{l-1}_{i,t}=\frac{1}{n}\sum f^{l-1}(\bm{i,t}),\forall\bm{(i,t)}\in D_{\text{edit}}\}$.
  • Transformations ($V$): The transformation refers to a specific layer in the LLM. During an edit, we fine-tune this transformation layer to encode the new knowledge. Each transformation $v(\cdot)$ associated with a specific key $\bm{k}$ stores the new weights with the updated knowledge. Typically, the transformation is fine-tuned with the model’s finetuning loss with updated knowledge.
@@
$v^{*}=\arg\min_{v}\bm{L}(f_{new}(i,t),y^{n}),$ (2)
where $\bm{L}$ is the language model loss, specifically the next-token prediction loss used by the base LLM. This loss ensures that the updated model $f_{\text{new}}$, when processing input $(i,t)$, generates the desired new answer $y_{n}$. Specifically, if the key is empty or the new fact falls outside the influence scope of existing keys, the transformation is directly finetuned from the original transformation layer. However, there may be instances where the new fact overlaps with the existing keys. In such cases, we finetune the transformation layer from the previously edited transformation to prevent catastrophic forgetting. Additionally, if the new key directly conflicts with previous edits, we will discard the previous entry and add a new one to update the knowledge. To ensure the universality, we primarily utilize the basic full fine-tuning approach as the transformation method. This involves adjusting the weights of the neural network to better align with the newly introduced or modified knowledge without altering the overall architecture of the model. The parameters that are tuned include all the weights within the specific layer of the network.
@@ 4.1 Datasets and Backbone Models
Backbone Models. Following previous work (Zheng et al., 2023), we adopt two vision language models as the base models. MiniGPT-4 (Zhu et al., 2023) is a powerful vision language model, leveraging Vicunna (Chiang et al., 2023) as the language model and a Vit-G/14 from EVA-CLIP (Sun et al., 2023) and a Q-former as the image encoder. BLIP-2 OPT (Li et al., 2023) utilizes a lightweight Q-former to bridge the gap between vision modality and text modality, where the ViT-L is adopted in the vision block, and the unsupervised-trained OPT model (Zhang et al., 2022) is used for decoder-based LLM.
Metrics. Following previous work (Zheng et al., 2023), we adopt the Editing Success Accuracy (Acc); Text Generality (T-Gen); Image Generality (I-Gen); and Locality (Loc) as the main metrics. To quantify the trade-off between generality and locality, we introduce the harmonic mean (HM) of the T-Gen, I-Gen and Loc. The detailed informations are in Appendix C.
@@
4.2 Baselines
4.3 Implementation Detail
In our comparisons of Finetuning, MEND and GRACE, we explore learning rates of $1.0$, $1e^{-1}$, $1e^{-2}$, $1e^{-3}$, $1e^{-4}$, and $1e^{-5}$. We observe that Finetuning, Memory, and MEND perform best with $1e^{-2}$. The choice of layer to edit is another hyperparameter for all editors. In all our editor comparisons, each editor modifies the same layer. For miniGPT-4, this is the dense layer of the llama block (llama_model.model.layers[31].mlp.up_proj), for BLIP-2 OPT moder, it is the OPT decoder layer (opt_model.model.decoder.layers[31].fc2). Recent work supporting the importance of selecting the correct layers to fine-tune corroborates this (Cheng et al., 2023). However, it’s important to note that the choice of layer is a practical hyperparameter: for comparison purposes, we ensure editors are compared when editing the same layers. For the distance function, we use the Euclidean distance if it is not explicitly mentioned. We select $\alpha$ using a small held-out set of only 5 unrelated samples. Since different models may have different latent feature distributions, $\alpha$ is treated as model-dependent. Once chosen, the same $\alpha$ is fixed per model (e.g., $\alpha$=0.2 for MiniGPT-4) for all evaluations.
4.4 Comparisons to Existing Methods
@@ 4.6 Efficiency Evaluation
Table 5: Time efficiency evaluation results on BLIP-2 OPT.
Time Efficiency. Training time is divided into two components, as detailed in Table 5. The first component is pre-training time, which involves either pre-training the model editing method or preparing the augmented index, such as training a meta-net. For instance, MEND, a meta-learning method, requires 22 hours to pre-train on 6,346 training samples. IKE, a retrieval-augmented in-context learning method, needs 12 hours to index 6,346 knowledge facts in advance.
@@ 4.8 Ablation Study
In this ablation study, we analyze the key design choices of BalancEdit. We first examine the effect of the hyperparameter $\alpha$, which controls the trade-off between generality and locality. We then evaluate the impact of alternative negative anchors and different distance functions to validate the generalizability. Additionally, we compare different negative sampling approaches and editing layers to assess the robustness of our method across different configurations. Finally, we conduct an extreme case analysis to demonstrate the behavior of BalancEdit in challenging scenarios.
@@
Effect of the Hyperparameter. In this study, we conducte a series of experiments on a subset of the OKVQA dataset to investigate how the parameter $\alpha$ affects the trade-off between generality and locality in model editing. As illustrated in Figure 5, we vary $\alpha$ from 0.1 and 0.3. The results, depicted in the figure, show that the editing success accuracy and text generality metrics consistently maintain a 100% accuracy rate. This stability is attributed to these metrics being closely tied to the key, with changes in the radius having no significant impact on them. However, the image generality metric, which is more challenging, shows a decline as $\alpha$ increases. This trend is anticipated because questions related to image generality tend to deviate from the key, despite sharing similar semantic content. Consequently, as the radius decreases, the edited model tends to overlook these questions. Conversely, the model’s performance on locality improves with an increase in $\alpha$. A smaller radius helps preserve the integrity of unrelated questions, ensuring that their answers remain unchanged. In this scenario, we observe that the harmonic mean of generality and locality initially increases and then decreases, further validating the existence of this trade-off. However, our method continues to achieve relatively high performance.
@@
Effect of Alternative Negative Anchors. To further validate the effectiveness of our approach, we conducte experiments using various negative anchors, including white negative images, on a subset of the OKEDIT dataset. As shown in Table 7, both black and white negative samples achieve 100% editing accuracy and exhibited a high harmonic mean in the locality-generality trade-off. The performance metrics for both white and black negative anchors, such as accuracy, generalization metrics, and locality, are remarkably consistent. The slight variations in the locality and I-Gen metrics suggest that white images can function as effective negative anchors, which also lack significant discriminative information. This consistency across different negative anchors highlights the robustness and adaptability of our pipeline in various settings and confirms our assumptions.
Method Edit Acc T-Generality I-Generality Locality HM Model
BalancEdit 100.00 98.89 65.38 61.18 71.85 BLIP-2 OPT
Random Negative Sample 100.00 100.00 49.12 65.08 65.61 BLIP-2 OPT
BalancEdit 100.00 99.87 76.46 53.14 71.58 MiniGPT-4
Random Negative Sample 100.00 99.00 66.93 45.92 64.08 MiniGPT-4
Table 8: Ablation study on negative sampling approaches
Effect of Negative Sampling Approaches. To verify our negative samples constuction approach, we compare our method with random negative sampling approach by randomly choosing irrelevant text-pair samples. As shown in Table. 8, BalancEdit consistently outperforms the random baseline in harmonic mean, demonstrating a better balance between generality and locality. This supports the effectiveness of our black image-based negative anchor, which offers a fact-agnostic, consistent, and efficient way to define a lower bound in the representation space. In contrast, random negative samples rely on external, unrelated examples and are often unstable in both quality and relevance, requiring assumptions that may not hold across diverse domains. Our method avoids these issues, making it more robust and scalable for real-world editing scenarios.
Effect of different editing layers.
Layer Acc T-Gen I-Gen Loc HM Model
30 100.00 100.00 64.03 66.39 73.75 MiniGPT4
31 100.00 99.87 76.46 53.14 71.58 MiniGPT4
30 100.00 100.00 78.43 44.99 66.70 BLIP-2 OPT
31 100.00 98.89 65.38 61.18 71.85 BLIP-2 OPT
Table 9: Ablation study on different editing layers.
To evaluate the robustness of our method across different editing layers, we apply our method to two separate layers, as shown in Table 9. The results show that our method maintains strong performance even when a different layer is chosen. Specifically, we observe higher harmonic mean performance on the MiniGPT-4 model, further demonstrating the robustness of our approach.
Extreme case analysation.
$\alpha$ Acc T-Gen I-Gen Loc HM
0 100.00 100.00 95.19 16.30 36.65
1 100.00 45.94 24.22 100.00 41.06
Table 10: Results of extreme cases of $\alpha$
In order to study the extreme case of hyperparameter, we conduct ablation experiments using extreme values of $\alpha$ on MiniGPT4 with OKEDIT dataset in Table 10. When $\alpha$ = 0, the influence radius is entirely determined by the negative sample, which leads to over-generalization and reduced locality, e.g., 16.30 locality score, as the edit is applied too broadly across the representation space. In contrast, when $\alpha$ = 1, the radius is determined solely by the positive sample, resulting in over-localization and poor generalization, e.g., 24.22 image generality score, since the edit is confined to a region that is too narrow. These results highlight the importance of balancing positive and negative influences to achieve both generality and locality.